OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry for the function? y = x2 -4x + 8 a: y = 2x b: x = 2 c: y = -2 d: x = -4

Mehek (mehek14):

what do you think?

OpenStudy (anonymous):

c

Mehek (mehek14):

the formula is \(\dfrac{-b}{2a}\)

Mehek (mehek14):

in your function, a = 1 , b = -4 \(\dfrac{-(-4)}{2*1}=\dfrac{4}{2}=2\)

Mehek (mehek14):

so B

OpenStudy (anonymous):

ok thanks

Mehek (mehek14):

yw